OpenStudy (anonymous):

Solve by factoring: 6x2 + x = 5 A. {1, 5} B. {-1, 5/6} C. {-1, 5} D. {1, 5/6}

OpenStudy (y2o2):

B. {1, 5/6} 6x²+x-5 = 0 (6x - 5)(x + 1)=0 x = 5/6 or x = -1

OpenStudy (anonymous):